A firm's product sells for $4 per unit in a highly competitive market. The firm produces output using capital (which it rents at $25 per hour) and labor (which is paid a wage of $30 per hour under a contract for 20 hours of labor services). Complete

the following table and use that information to answer these questions. (LOI, LO2,

LO5, LO6)


a. Identify the fixed and variable inputs.

b. What are the firm's fixed costs?

c. What is the variable cost of producing 475 units of output?

d. How many units of the variable input should be used to maximize profits?

e. What are the maximum profits this firm can earn?

f. Over what range of the variable input usage do increasing marginal returns exist?

8. Over what range of the variable input usage do decreasing marginal returns exist?

h. Over what range of input usage do negative marginal returns exist?


1
Expert's answer
2022-03-10T17:58:50-0500

a). In this case labor is a fixed input while capital is a variable input

b). The fixed cost is determined by multiplying the labor cost per hour and the number of hours. That is $30*20 = $600

c). Variable cost of producing 475 units is equal to units of capital at 475th unit multiplied by the rental fee. (7*25) = $175

d). A firm maximizes production when the marginal product is equal to the price of the variable input. therefore, at point of maximization MPK = rental fee. Since rent is $25, then the maximizing point is when the firm produces 6 units.

e) Profit = total revenue - total cost

Total revenue = 475*4 = $1900

Total cost = Fixed cost + variable cost = 600+ (25*6) = 750

profit = 1900-750 = $1150

f) The point of increasing marginal return is between 1 unit to 4 units of capital.

g). The point of decreasing marginal return is between 5 units to 7 units of capital

h). The point of negative marginal return is 8 units and above.